The length of a rectangle is 3 in. more than its width. The perimeter is 30 in. Write and solve the equation to find the width of the rectangle. Formula is P=2 x length + 2 x width
Please show all work.
Thanks

Answer by nerdybill(7384)

Let w = width of rectangle
then because "length of a rectangle is 3 in. more than its width" we have
w+3 = length of rectangle
.
Plugging all the information into:
P=2 x length + 2 x width
we get
30 = 2(w+3) + 2x
dividing both sides by 2:
15 = w+ 3 + x
15 = 2w + 3
12 = 2w
6 inches = w (width)